OpenStudy (anonymous):

Its a 7th grade ques, very easy. Please help. Construct an equation from the information given and then solve it : 9 times a number has 1 subtracted from it. The result is trebled. The answer is the same if you add 7 to twice the number and multiply the result by 7. What is the number ?

OpenStudy (anonymous):

plz help directix

OpenStudy (anonymous):

directix are u there

OpenStudy (anonymous):

shall i give it a shot ?

OpenStudy (anonymous):

3(9x-1) = 7(2x+7)

Directrix (directrix):

Let x be the number. 9x - 1 is 9 times the number with one subtracted from it. The result is trebled. (means multiplied by 3) 3*(9x - 1) The answer is the same if you add 7 to twice the number and multiply the result by 7. 7*(2x + 7) So, solve this: 3*(9x - 1) = 7*(2x + 7)

OpenStudy (anonymous):

so i was right

Directrix (directrix):

Right. What did you get for the number?

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

27x - 3 = 14x + 49

OpenStudy (anonymous):

27x-14x=52

OpenStudy (anonymous):

13x=52

OpenStudy (anonymous):

x=52/13 x=4

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

am i right ?

Directrix (directrix):

That is what I got.

OpenStudy (anonymous):

thanks

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

i'll take a leave. bye

Directrix (directrix):

Me. too.